What is the answer to this problem when you multiply
7w(2w - 3)(2w + 3) What is the process for solving this problem?

Thank you.

Answer by malakumar_kos@yahoo.com(315)


What is the answer to this problem when you multiply
7w(2w - 3)(2w + 3) What is the process for solving this problem

7w(2w+3)(2w-3) = 7w[(2w)^2-(3)^2] use (a+b)(a-b) = a^2-b^2
= 7w[4w^2-9]
= 7w*4w^2-7w*9
= 28w^3-63w is the required product